Lisp-Stat Information

Luke Tierney School of Statistics University of Minnesota

September 30, 1998

Lisp-Stat [4] is an extensible statistical computing environment for data analysis, statistical instruction and research, with an emphasis on providing a framework for exploring the use of dynamic graphical methods. Extensibility is achieved by basing Lisp-Stat on the Lisp language, in particular on a subset of Common Lisp. Lisp-Stat extends standard Lisp arithmetic operations to perform element-wise operations on lists and vectors, and adds a variety of basic statistical and linear algebra functions. A portable window system interface forms the basis of a dynamic graphics system that is designed to work identically in a number of different graphical user interface environments, such as the Macintosh operating system, the X window system, and Microsoft Windows. A prototype-based object-oriented programming system is used to implement the graphics system and to allow it to be customized and adapted. The object-oriented programming system is also used as the basis for statistical model representations, such as linear and nonlinear regression models and generalized linear models [4, 5]. Many aspects of the system design were motivated by the S language [1].

1 Obtaining Lisp-Stat

Implementations of XLISP-STAT, a version of Lisp-Stat based on the XLISP language, are available for the Macintosh, UNIX systems running X11, and for MS Windows (16bit versions and 32bit versions are available).

The most recent versions of xlispstat is available by ftp from several sites. The primary ftp site is

ftp.stat.umn.edu/pub/xlispstat. This site is mirrorred at

• StatLib^{1}

• UCLA^{2}

• Perugia, Italy^{3}, also accessible as an html document^{4}

• Heidelberg, Germany^{5}

• Vienna, Austria^{6}

• Sydney, Australia^{7}

More recent development snapshots may be available in the snapshots directory^{8}.

UNIX versions are currently available only in source form. Binary version for standard configurations may be made available in the near future.

1http://www.stat.cmu.edu

2http://www.stat.ucla.edu

3ftp://ftp.stat.unipg.it/pub/stat/xlispstat

4http://www.stat.unipg.it/IASC/Misc-stat-soft-XLS.html

5ftp://statlab.uni-heidelberg.de/pub/mirrors/umn

6http://sunsite.univie.ac.at/

7ftp://srsuna.shlrc.mq.edu.au/pub/xlispstat

8http://www.stat.umn.edu/ luke/xls/projects/projects.html

2 Lisp-Stat Documentation

There primary documentation is the book LISP-STAT: An Object-Oriented Environment for Statistical Computing and Dynamic Graphics [4]. Additional documentation:

• A tutorial introduction^{9} [3] corresponding roughly to the second chapter in the book and some of the book’s appendices.

• A technical report^{10} describing the generalized linear models system now included with the distribution [5].

• Notes^{11} on the MS Windows version [8].

• Notes^{12} on changes from the version documented in the book [7].

Please let me know if you have other items you would like me to list.

3 Projects Using Lisp-Stat

Several larger projects have been implemented using Lisp-Stat as a base. Here are a few examples. Please let me know if there are others you would like me to list.

• Rcode, the code to accompany Cook and Weisberg [2].

• ViSta [9]

• KDE: Lisp-Stat objects for Kernel Density Estimation

• Steven D. Majewski’s Xray Energy Spectra analysis package written in LispStat.

The UCLA archive contains a variety of useful contributed code.

4 Future Directions

Lisp-Stat was first release in 1989. It has been used for data analysis, as a research tool, and for implementing several larger projects (e. g. [2]; [9]). Based on experience gained from this use, the system is currently being redesigned. The redesign is evolutionary, with backward compatibility a major objective. The redesign project can be divided into six major segments: the basic Lisp system, data representation and operating system issues, the object system, the graphical system, the statistical component, and the user interface. Some more details are given in Tierney [6].

A projects page lists recent developments and work in progress.

A talk on “New Developments in Lisp-Stat” was presented at the 1998 Interface meeting in Minneapolis. Slides are available as a gziped postscript file. The proceedings paper is available as an html document and as a gziped postscript file.

5 Further Information

The internet mailing list stat-lisp-news is available for discussion of issues related to statistical computing using Lisp. To join the list, send a message to majordomo@stat.umn.edu containing the line

        subscribe stat-lisp-news

9/ luke/xls/tutorial/techreport/techreport.html

10/ luke/xls/glim/glim/glim.html

11/ luke/xls/wxls/wxls/wxls.html

12/ luke/xls/changes/changes/changes.html

or

    subscribe stat-lisp-news ADDRESS

with ADDRESS replaced by the email address you want to subscribe.

An archive of the mailing list is available. A second mailing list for issues related to Lisp-Stat development has been established recently. The list is

called stat-lisp-devel. It is also managed by majordomo@stat.umn.edu; an archive of this list is available as well.

References

[1] Becker, R. A., Chambers, J. M., and Wilks, A. R. (1988), The new S language: A Programming Environment for Data Analysis and Graphics, Pacific Grove, CA: Wadsworth.

[2] Cook, R. D. and Weisberg, S. (1994), An Introduction to Regression Graphics, New York, NY: Wiley.

URL: http://www.stat.umn.edu/~rcode .

[3] Tierney, L. (1989), “XLISP-STAT: A Statistical Environment Based on the XLISP Language (Version 2.0),” Technical Report No. 528, School of Statistics, University of Minnesota.

URL (html): http://www.stat.umn.edu/~luke/xls/tutorial/techreport/techreport.html .
URL (PostScript): ftp://ftp.stat.umn.edu/pub/xlispstat/doc/techreport.ps.

[4] Tierney, L. (1990), LISP-STAT: An Object-Oriented Environment for Statistical Computing and Dynamic Graphics, NewYork, NY: Wiley. ISBN: 0-471-50916-7

[5] Tierney, L. (1991), “Generalized Linear Models in Lisp-Stat,” Technical Report No. 557, School of Statistics, University of Minnesota.

URL (html): http://www.stat.umn.edu/~luke/xls/glim/glim/glim.html .
URL (PostScript): ftp://ftp.stat.umn.edu/pub/xlispstat/doc/glim.ps.

[6] Tierney, L. (1995), “Recent developments and future directions in Lisp-Stat,” Technical Report No. 608, School of Statistics, University of Minnesota.

URL: http://www.stat.umn.edu/PAPERS/tr608.html.

[7] Tierney, L., “XLISP-STAT 2.1 Release 3 Beta Release Notes.”

URL (html): http://www.stat.umn.edu/~luke/xls/changes/changes/changes.html .
URL (PostScript): ftp://ftp.stat.umn.edu/pub/xlispstat/doc/changes.ps.

[8] Tierney, L., “XLISP-STAT for Microsoft Windows 3.1.”

URL (html): http://www.stat.umn.edu/~luke/xls/wxls/wxls/wxls.html .
URL (PostScript): ftp://ftp.stat.umn.edu/pub/xlispstat/doc/wxls.ps.

[9] Young, F. W. (1993), “ViSta – The visual statistics system. A research and development test bed for statistical visualization techniques,” Technical Report, Psychometrics Lab, University of North Carolina.

URL: http://forrest.psych.unc.edu/research/ViSta.html.